# DSCA: A Digital Subtraction Angiography Sequence Dataset and Spatio-Temporal Model for Cerebral Artery Segmentation

### v1

@@ III-B TemporalFormer
The dynamic flows within DSA sequences possess rich temporal information, characterized by meaningful contextual correlations among frames. Consequently, this temporal information offers valuable assistance in segmenting cerebral arteries and categorizing them into BV and MTA. To effectively capture the global context and inter-frame correlations, we feed the features generated by each frame in the TEB into the TemporalFormer (TF), as depicted in Fig. 3.
We consider the features $F_{s}\in\mathbb{R}^{N\times c\times h\times w}$ extracted from the TEB, where each frame feature is defined as $f^{t}_{s}\in\mathbb{R}^{B\times c\times h\times w}$. Here, $c$, $h$, and $w$ respectively represent the number of channels, the height, and the width of $F_{s}$. The index $t\in\{0,T\}$ is treated as a token and inputted into the TF. Specifically, $F_{s}$ is initially reshaped into $\in\mathbb{R}^{B\times T\times c\times h\times w}$ to restore its original dimensions. Then, it is converted to a feature vector $\mathcal{F}_{s}$ $\in\mathbb{R}^{(Bhw)\times T\times c}$ by flattening $B$, $h$, and $w$. Subsequently, we integrate learnable position encoding $f_{p}\in\mathbb{R}^{T\times c}$ with $\mathcal{F}_{s}$ by adding them together. Following this, $\mathcal{F}_{s}$ is transformed into $\in\mathbb{R}^{B\times(hwT)\times c}$ to generate a residual feature denoted as $\mathcal{F}^{\prime}_{s}$, which is then restored. Afterwards, $\mathcal{F}_{s}$ undergoes a temporal multi-head self-attention operation, which can be formulated as:
@@
Figure 4: The details of the dimension transformation process in TemporalFormer.
@@ III-C Spatio-Temporal Fusion Module
To effectively fuse the spatial and temporal features from the encoder, we introduce an STF module. Given the feature $F_{m}\in\mathbb{R}^{B\times c\times h\times w}$ from the SEB and the feature $F_{s}\in\mathbb{R}^{B\times c\times h\times w}$ from the TF module, we initially apply Maxpool to both $F_{m}$ and $F_{s}$ to reduce the size of feature maps, yielding $F^{\prime}_{m}$, and $F^{\prime}_{s}$ $\in\mathbb{R}^{B\times c\times\frac{h}{2}\times\frac{w}{2}}$. Subsequently, the two features are passed through a shared convolutional block to generate feature vector $Q_{m},K_{m},V_{m},Q_{s},K_{s},V_{s}$ $\in\mathbb{R}^{B\times c\times\frac{hw}{4}}$. Afterwards, $Q_{m},Q_{s}$ are multiplied by $K_{m}^{\mathtt{T}},K_{s}^{\mathtt{T}}$, respectively, and then dividing by $\sqrt{c}$. The weight matrices $\alpha_{i}$, $\alpha_{s}$ $\in\mathbb{R}^{B\times\frac{hw}{4}}$ are obtained through the softmax function, similar to Equation (2). The two weight matrices are then summed, and the result is respectively multiplied by $V_{m}$ and $V_{s}$ to get $\tilde{F}^{\prime}_{m}$, $\tilde{F}^{\prime}_{s}$ $\in\mathbb{R}^{B\times\frac{hw}{4}}$, as depicted below:
@@ III-D Loss Function
TABLE II: Quantitative comparisons with DSA segmentation and general segmentation methods on DSCA for BV and MTA. The results are average±$\pm$±standard deviation obtained from the five-fold cross-validation. All metrics are expressed as percentages (%).
Methods JAC Dice SEN PRE AUC
BV MTA BV MTA BV MTA BV MTA BV MTA
DSA Zhang et al. [26] 67.56±$\pm{0.82}$ 67.45±$\pm{1.36}$ 80.11±$\pm{0.61}$ 73.00±$\pm{1.26}$ 76.70±$\pm{0.67}$ 71.74±$\pm{1.46}$ 84.61±$\pm{0.77}$ 75.9±$\pm{1.09}$ 88.08±$\pm{0.34}$ 77.20±$\pm{1.31}$
MDCNN [7] 66.16±$\pm{2.95}$ 69.57±$\pm{0.64}$ 78.95±$\pm{2.23}$ 74.61±$\pm{0.43}$ 74.05±$\pm{4.27}$ 74.12±$\pm{1.56}$ 85.69±$\pm{1.15}$ 77.09±$\pm{1.38}$ 86.79±$\pm{2.10}$ 79.06±$\pm{0.77}$
ERNet [27] 73.39±$\pm{0.67}$ 78.01±$\pm{1.07}$ 84.19±$\pm{0.46}$ 82.61±$\pm{1.18}$ 81.67±$\pm{1.56}$ 80.75±$\pm{1.50}$ 87.49±$\pm{1.01}$ 87.80±$\pm{2.02}$ 90.60±$\pm{0.76}$ 86.03±$\pm{1.54}$
General UNet [8] 71.30±$\pm{0.04}$ 73.24±$\pm{1.76}$ 82.74±$\pm{0.33}$ 77.68±$\pm{1.67}$ 80.10±$\pm{0.39}$ 77.24±$\pm{1.69}$ 86.24±$\pm{0.30}$ 80.18±$\pm{2.59}$ 89.80±$\pm{0.19}$ 81.77±$\pm{1.62}$
CENet [33] 68.34±$\pm{0.65}$ 79.97±$\pm{2.59}$ 80.69±$\pm{0.51}$ 84.32±$\pm{2.78}$ 76.86±$\pm{2.04}$ 82.85±$\pm{2.73}$ 85.78±$\pm{2.16}$ 87.20±$\pm{2.67}$ 88.18±$\pm{0.97}$ 87.09±$\pm{2.52}$
CSNet [16] 73.02±$\pm{0.93}$ 81.28±$\pm{1.87}$ 83.95±$\pm{0.65}$ 85.37±$\pm{1.79}$ 80.77±$\pm{2.05}$ 84.11±$\pm{2.07}$ 87.98±$\pm{1.23}$ 89.13±$\pm{1.63}$ 90.17±$\pm{0.99}$ 88.85±$\pm{2.24}$
AttUNet [18] 69.88±$\pm{0.78}$ 74.58±$\pm{1.00}$ 81.73±$\pm{0.57}$ 79.10±$\pm{0.92}$ 76.51±$\pm{1.59}$ 77.85±$\pm{1.26}$ 88.47±$\pm{1.12}$ 82.68±$\pm{1.01}$ 88.07±$\pm{0.77}$ 82.31±$\pm{1.13}$
SwinUNet [20] 64.22±$\pm{0.44}$ 67.41±$\pm{1.45}$ 77.61±$\pm{0.35}$ 73.45±$\pm{1.30}$ 75.33±$\pm{0.50}$ 71.92±$\pm{1.28}$ 79.92±$\pm{1.69}$ 75.79±$\pm{1.49}$ 87.33±$\pm{0.25}$ 77.28±$\pm{1.22}$
MISSFormer [34] 70.09±$\pm{0.37}$ 78.97±$\pm{3.67}$ 81.89±$\pm{0.21}$ 84.06±$\pm{3.37}$ 79.10±$\pm{0.62}$ 82.41±$\pm{3.39}$ 85.58±$\pm{0.91}$ 86.87±$\pm{3.54}$ 89.30±$\pm{0.30}$ 87.08±$\pm{3.19}$
H2Former [35] 69.47±$\pm{0.62}$ 83.90±$\pm{3.79}$ 81.47±$\pm{0.43}$ 88.34±$\pm{3.84}$ 77.85±$\pm{1.54}$ 86.87±$\pm{3.62}$ 86.21±$\pm{1.17}$ 92.51±$\pm{4.42}$ 88.67±$\pm{0.74}$ 91.59±$\pm{3.82}$
TransUNet [19] 72.43±$\pm{0.46}$ 81.38±$\pm{1.24}$ 83.63±$\pm{0.34}$ 85.09±$\pm{1.19}$ 80.99±$\pm{0.92}$ 83.89±$\pm{1.12}$ 87.20±$\pm{0.45}$ 88.33±$\pm{2.08}$ 90.24±$\pm{0.45}$ 88.29±$\pm{1.48}$
nnUNet [22] 76.76±$\pm{0.45}$ 82.54±$\pm{2.42}$ 86.45±$\pm{0.31}$ 86.22±$\pm{2.50}$ 85.29±$\pm{0.45}$ 85.70±$\pm{2.50}$ 88.06±$\pm{0.59}$ 88.08±$\pm{2.20}$ 92.41±$\pm{0.22}$ 89.87±$\pm{2.51}$
DSANet 78.09±$\mathbf{\pm{0.20}}$ 88.22±$\mathbf{\pm{1.52}}$ 87.32±$\mathbf{\pm{0.13}}$ 92.26±$\mathbf{\pm{1.56}}$ 86.40±$\mathbf{\pm{0.39}}$ 91.37±$\mathbf{\pm{1.57}}$ 88.59±$\mathbf{\pm{0.53}}$ 94.01±$\mathbf{\pm{1.40}}$ 92.98±$\mathbf{\pm{0.18}}$ 94.98±$\mathbf{\pm{1.39}}$
@@ IV-A Implementation Details
The preprocessing pipelines vary between training and testing. For training, we utilized a five-fold cross-validation method by dividing the 180 sequences in the training set, with four folds for training and one for validation. Each batch comprised 2 samples at 512$\times$512 pixels which were randomly cropped from images of varying dimensions. Additionally, common data augmentation methods were applied to optimize the training process, including random rotation, elastic deformation, random scaling, random cropping, gamma augmentation, and mirroring. For testing, we only perform mirroring operation for data augmentation. The testing procedure involved cropped patches of 512$\times$512 pixels from the test image via sliding windows. We applied padding with mirroring to ensure consistency in patch sizes. Subsequently, these patches were inputted into the network for prediction and then sequentially reconstructed to match the original image size. Finally, we compared the reconstructed predictions with the GT across the entire image. The performance metrics were determined based on the average values calculated by the five-fold cross-validation at the image level on the test set.
@@ IV-B Evaluation Metrics
To comprehensively evaluate the segmentation performance, the following metrics are calculated and compared: Jaccard (JAC) = TP / (TP + FP + FN); Dice = 2$\times$×TP / (2$\times$×TP + FP + FN); Sensitivity (SEN) = TP / (TP + FN); Precision (PRE) = TP / (TP + FP); Area Under the ROC Curve (AUC), where TP, FP, TN, and FN are true positive, false positive, true negative, and false negative, respectively. The clDice [36] is also used to evaluate vessel connectivity when measuring the entire CA.
Figure 5: Segmentation results on the DSCA dataset. Green is MTA, and red is BV. Enlarged viewing for better clarity.
IV-C Cerebrovascular Segmentation Results
We extensively assessed our proposed method on the DSCA dataset, comparing it with state-of-the-art segmentation approaches, which include UNet [8], CENet [33], CSNet [16], AttentionUNet [17], SwinUNet [20], MISSFormer [34], H2Former [35], TransUNet [19], nnUNet [22], and [27]. Additionally, we also reproduced [26] and [7] for comparison, as their codes are not publicly available. The comparative results of segmentation are shown in Table II.
We can observe that DSANet achieves the best performance across all metrics for segmenting both MTA and BV, which indicates the superiority of our method. In MTA segmentation, DSANet holds a significant advantage over other methods in terms of JAC, Dice, SEN, and AUC, surpassing H2Former by 4.32%, 3.92%, 4.50%, and 3.39%, respectively. A statistical significance of (p$<$0.05) was obtained using a paired t-test between the two methods. In BV segmentation, the pixels of BV constitute a relatively small proportion of the total image pixels, which explains why our proposed method is not significantly ahead of the comparative method. Nevertheless, DSANet stands out as the only method higher than 78% in JAC and 86% in SEN, surpassing nnUNet by 1.33% and 1.11%, respectively. Similarly, it also outperforms nnUNet by 0.87% in Dice and 0.57% in AUC, highlighting its superior performance with statistical significance (p$<$0.05).
For a more intuitive comparison of CA segmentation performance across different methods, we also provide a visual comparison in Fig. 5. As we can see the shape and background of the CA are very intricate, with low contrast in small vessels. In this challenging context, our DSANet effectively segments the CA, particularly ensuring small vessels with more robust connectivity, as indicated in the detailed orange boxes in ICA, ECA, and VA. In addition, skulls with deep contrast are mistakenly extracted as vessels in other methods, while our method successfully suppresses such false positives, as shown by the white arrows in ECA and VA in Fig. 5. In ICA, CENet and TransUNet perform well in segmenting vessels within the left orange box, while their performance diminishes in the right orange box. In contrast, DSANet achieves superior segmentation performance in both boxes. Furthermore, other methods erroneously segment the MTA in ECA (depicted in green), while DSANet effectively addresses this issue.
We also compare the performance of the entire CA without distinguishing BV and MTA to demonstrate the superiority of DSANet further. As indicated in Table III, DSANet holds a slight advantage over other methods in JAC, Dice, SEN, and AUC, surpassing the second-place method by 1.06%, 0.64%, 0.85%, and 0.43%, respectively, with statistical significance (p$<$0.05). To achieve precise segmentation, it is essential to focus on the intricate details of small vessel structures, which represent a minority of the total image pixels. Thus, clDice [36] is employed to assess vessel connectivity. DSANet significantly outperforms other methods, surpassing nnUNet by 1.04% in clDice with statistical significance (p$<$0.001).
@@ IV-C Cerebrovascular Segmentation Results
TABLE III: Quantitative comparisons with DSA segmentation and general segmentation methods on DSCA for the entire vessel. The results are average±$\pm$±standard deviation obtained from five-fold cross-validation. All metrics are based on percentages (%).
@@
TABLE IV: Ablation study of these components of the proposed DSANet. The results are average±$\pm$±standard deviation obtained from the five-fold cross-validation experiment. All metrics are expressed as percentages (%).

### v2

@@ III-B TemporalFormer
The dynamic flows within DSA sequences possess rich temporal information, characterized by meaningful contextual correlations among frames. Consequently, this temporal information offers valuable assistance in segmenting cerebral arteries and categorizing them into BV and MAT. To effectively capture the global context and inter-frame correlations, we feed the features generated by each frame in the TEB into the TemporalFormer (TF), as depicted in Fig. 3.
We consider the features $F_{s}\in\mathbb{R}^{(BT)\times c\times h\times w}$ extracted from the TEB, where each frame feature is defined as $f^{t}_{s}\in\mathbb{R}^{B\times c\times h\times w}$. Here, $c$, $h$, and $w$ respectively represent the number of channels, the height, and the width of $F_{s}$. The index $t\in\{0,T\}$ is treated as a token and inputted into the TF. Specifically, $F_{s}$ is initially reshaped into $\in\mathbb{R}^{B\times T\times c\times h\times w}$ to restore its original dimensions. Then, it is converted to a feature vector $\mathcal{F}_{s}$ $\in\mathbb{R}^{(Bhw)\times T\times c}$ by flattening $B$, $h$, and $w$. Subsequently, we integrate learnable position encoding $f_{p}\in\mathbb{R}^{T\times c}$ with $\mathcal{F}_{s}$ by adding them together. Following this, $\mathcal{F}_{s}$ is transformed into $\in\mathbb{R}^{B\times(hwT)\times c}$ to generate a residual feature denoted as $\mathcal{F}^{\prime}_{s}$, which is then restored. Afterwards, $\mathcal{F}_{s}$ undergoes a temporal multi-head self-attention operation, which can be formulated as:
@@
Figure 4: The details of the dimension transformation process in TemporalFormer, where c is 512.
@@ III-C Spatio-Temporal Fusion Module
To effectively fuse the spatial and temporal features from the encoder, we introduce an STF module. Given the feature $F_{m}\in\mathbb{R}^{B\times c\times h\times w}$ from the SEB and the feature $F_{s}\in\mathbb{R}^{B\times c\times h\times w}$ from the TF module, we initially apply Maxpool to both $F_{m}$ and $F_{s}$ to reduce the size of feature maps and reduce noise interference to retain critical features, yielding $F^{\prime}_{m}$, and $F^{\prime}_{s}$ $\in\mathbb{R}^{B\times c\times\frac{h}{2}\times\frac{w}{2}}$. Subsequently, the two features are passed through a shared convolutional block to generate feature vector $Q_{m},K_{m},V_{m},Q_{s},K_{s},V_{s}$ $\in\mathbb{R}^{B\times c\times\frac{hw}{4}}$. Afterward, $Q_{m},Q_{s}$ are multiplied by $K_{m}^{\mathtt{T}},K_{s}^{\mathtt{T}}$, respectively, and then dividing by $\sqrt{c}$ to prevent the dot product from becoming too large and causing the vanishing gradient problem. The weight matrices $\alpha_{i}$, $\alpha_{s}$ $\in\mathbb{R}^{B\times\frac{hw}{4}}$ are obtained through the softmax function, similar to Equation (2). Additionally, we believe that spatio-temporal features at the same position can enhance the representation of the network. Therefore, the two weight matrices are then summed, and the result is respectively multiplied by $V_{m}$ and $V_{s}$ to get the enhanced features $\tilde{F}^{\prime}_{m}$, $\tilde{F}^{\prime}_{s}$ $\in\mathbb{R}^{B\times\frac{hw}{4}}$, as depicted below:
@@ IV-A Implementation Details
The preprocessing pipelines vary between training and testing. For training, we utilized a five-fold cross-validation method by dividing the 180 sequences in the training set, with four folds for training and one for validation. Each batch comprised 2 samples at 512$\times$512 pixels which were randomly cropped from images of varying dimensions. Both our method and the comparison methods used data pre-processed by nnUNet, and no post-processing was applied as it may result in the removal of small vessels. This post-processing method, which discards fractured small vessels, is unreasonable and may adversely affect the performance of comparative experiments, especially the clDice [41] metric. Additionally, common data augmentation methods were applied to optimize the training process, including random rotation, elastic deformation, random scaling, random cropping, gamma augmentation, and mirroring. For testing, we only perform mirroring operations for data augmentation. The testing procedure involved cropped patches of 512$\times$512 pixels from the test image via sliding windows. We applied padding with mirroring to ensure consistency in patch sizes. Subsequently, these patches were inputted into the network for prediction and then sequentially reconstructed to match the original image size. Finally, we compared the reconstructed predictions with the GT across the entire image. The performance metrics on the test set were obtained by averaging the quantitative segmentation results obtained from the five models based on the five-fold cross-validation.
@@ IV-B Evaluation Metrics
To comprehensively evaluate the segmentation performance, the following metrics are calculated and compared: Jaccard (JAC)=TP/(TP+FP+FN); Dice=2$\times$×TP/(2$\times$×TP+FP+FN); Sensitivity (SEN)=TP/(TP+FN); Precision (PRE)=TP/(TP+FP); Area Under the ROC Curve (AUC), where TP, FP, TN, and FN are true positive, false positive, true negative, and false negative, respectively. The clDice [41] is also used to evaluate vessel connectivity when measuring the entire cerebral artery. Moreover, a paired t-test is conducted to calculate the p-values.
@@
TABLE III: Quantitative comparisons with DSA segmentation and general segmentation methods on DSCA for BV and MAT. The results are average±$\pm$±standard deviation obtained from the five-fold cross-validation. All metrics are expressed as percentages (%).
Methods JAC Dice SEN PRE AUC
BV MAT BV MAT BV MAT BV MAT BV MAT
DSA Zhang et al. [28] 67.56 $\pm$ 0.82 67.45 $\pm$ 1.36 80.11 $\pm$ 0.61 73.00 $\pm$ 1.26 76.70 $\pm$ 0.67 71.74 $\pm$ 1.46 84.61 $\pm$ 0.77 75.9 $\pm$ 1.09 88.08 $\pm$ 0.34 77.20 $\pm$ 1.31
MDCNN [7] 66.16 $\pm$ 2.95 69.57 $\pm$ 0.64 78.95 $\pm$ 2.23 74.61 $\pm$ 0.43 74.05 $\pm$ 4.27 74.12 $\pm$ 1.56 85.69 $\pm$ 1.15 77.09 $\pm$ 1.38 86.79 $\pm$ 2.10 79.06 $\pm$ 0.77
ERNet [31] 73.39 $\pm$ 0.67 78.01 $\pm$ 1.07 84.19 $\pm$ 0.46 82.61 $\pm$ 1.18 81.67 $\pm$ 1.56 80.75 $\pm$ 1.50 87.49 $\pm$ 1.01 87.80 $\pm$ 2.02 90.60 $\pm$ 0.76 86.03 $\pm$ 1.54
General UNet [8] 71.30 $\pm$ 0.04 73.24 $\pm$ 1.76 82.74 $\pm$ 0.33 77.68 $\pm$ 1.67 80.10 $\pm$ 0.39 77.24 $\pm$ 1.69 86.24 $\pm$ 0.30 80.18 $\pm$ 2.59 89.80 $\pm$ 0.19 81.77 $\pm$ 1.62
CENet [42] 68.34 $\pm$ 0.65 79.97 $\pm$ 2.59 80.69 $\pm$ 0.51 84.32 $\pm$ 2.78 76.86 $\pm$ 2.04 82.85 $\pm$ 2.73 85.78 $\pm$ 2.16 87.20 $\pm$ 2.67 88.18 $\pm$ 0.97 87.09 $\pm$ 2.52
CSNet [16] 73.02 $\pm$ 0.93 81.28 $\pm$ 1.87 83.95 $\pm$ 0.65 85.37 $\pm$ 1.79 80.77 $\pm$ 2.05 84.11 $\pm$ 2.07 87.98 $\pm$ 1.23 89.13 $\pm$ 1.63 90.17 $\pm$ 0.99 88.85 $\pm$ 2.24
AttUNet [18] 69.88 $\pm$ 0.78 74.58 $\pm$ 1.00 81.73 $\pm$ 0.57 79.10 $\pm$ 0.92 76.51 $\pm$ 1.59 77.85 $\pm$ 1.26 88.47 $\pm$ 1.12 82.68 $\pm$ 1.01 88.07 $\pm$ 0.77 82.31 $\pm$ 1.13
SwinUNet [20] 64.22 $\pm$ 0.44 67.41 $\pm$ 1.45 77.61 $\pm$ 0.35 73.45 $\pm$ 1.30 75.33 $\pm$ 0.50 71.92 $\pm$ 1.28 79.92 $\pm$ 1.69 75.79 $\pm$ 1.49 87.33 $\pm$ 0.25 77.28 $\pm$ 1.22
MISSFormer [43] 70.09 $\pm$ 0.37 78.97 $\pm$ 3.67 81.89 $\pm$ 0.21 84.06 $\pm$ 3.37 79.10 $\pm$ 0.62 82.41 $\pm$ 3.39 85.58 $\pm$ 0.91 86.87 $\pm$ 3.54 89.30 $\pm$ 0.30 87.08 $\pm$ 3.19
H2Former [44] 69.47 $\pm$ 0.62 83.90 $\pm$ 3.79 81.47 $\pm$ 0.43 88.34 $\pm$ 3.84 77.85 $\pm$ 1.54 86.87 $\pm$ 3.62 86.21 $\pm$ 1.17 92.51 $\pm$ 4.42 88.67 $\pm$ 0.74 91.59 $\pm$ 3.82
TransUNet [19] 72.43 $\pm$ 0.46 81.38 $\pm$ 1.24 83.63 $\pm$ 0.34 85.09 $\pm$ 1.19 80.99 $\pm$ 0.92 83.89 $\pm$ 1.12 87.20 $\pm$ 0.45 88.33 $\pm$ 2.08 90.24 $\pm$ 0.45 88.29 $\pm$ 1.48
nnUNet [22] 76.76 $\pm$ 0.45 82.54 $\pm$ 2.42 86.45 $\pm$ 0.31 86.22 $\pm$ 2.50 85.29 $\pm$ 0.45 85.70 $\pm$ 2.50 88.06 $\pm$ 0.59 88.08 $\pm$ 2.20 92.41 $\pm$ 0.22 89.87 $\pm$ 2.51
Sequence + MinIP nnUNet* [22] 77.52 $\pm$ 0.22 85.90 $\pm$ 2.10 86.94 $\pm$ 0.17 89.88 $\pm$ 2.29 85.73 $\pm$ 0.32 88.76 $\pm$ 2.09 88.56 $\pm$ 0.31 91.93 $\pm$ 2.57 92.64 $\pm$ 0.15 93.00 $\pm$ 1.88
UNet* [8] 62.12 $\pm$ 1.17 56.67 $\pm$ 5.80 75.97 $\pm$ 0.86 64.74 $\pm$ 4.48 84.39 $\pm$ 2.53 89.24 $\pm$ 6.32 70.78 $\pm$ 2.87 75.26 $\pm$ 2.21 91.48 $\pm$ 1.16 70.50 $\pm$ 3.16
MISSFormer* [43] 65.55 $\pm$ 1.98 70.69 $\pm$ 3.53 78.69 $\pm$ 1.49 77.09 $\pm$ 3.18 87.95 $\pm$ 0.69 73.89 $\pm$ 3.99 72.28 $\pm$ 2.68 83.34 $\pm$ 4.49 93.29 $\pm$ 0.25 82.82 $\pm$ 3.82
H2Former* [44] 63.55 $\pm$ 1.13 60.23 $\pm$ 5.53 77.14 $\pm$ 0.87 67.91 $\pm$ 5.28 84.68 $\pm$ 0.39 61.60 $\pm$ 5.77 72.05 $\pm$ 1.38 87.98 $\pm$ 1.34 91.66 $\pm$ 0.18 78.51 $\pm$ 2.98
DeepMedic [30] 60.37 $\pm$ 1.42 58.76 $\pm$ 1.68 74.46 $\pm$ 1.01 67.76 $\pm$ 1.02 72.08 $\pm$ 3.35 74.48 $\pm$ 1.93 78.63 $\pm$ 3.60 63.49 $\pm$ 3.26 85.66 $\pm$ 1.57 77.99 $\pm$ 0.93
DSNet [34] 76.62 $\pm$ 0.32 78.82 $\pm$ 2.59 86.32 $\pm$ 0.23 82.76 $\pm$ 2.68 84.85 $\pm$ 0.53 81.82 $\pm$ 2.68 88.23 $\pm$ 3.25 85.25 $\pm$ 3.30 92.20 $\pm$ 2.59 85.88 $\pm$ 2.83
CAVE [45] 69.11 $\pm$ 2.80 79.09 $\pm$ 2.16 81.12 $\pm$ 1.88 83.70 $\pm$ 2.15 83.26 $\pm$ 4.51 81.90 $\pm$ 2.28 80.48 $\pm$ 1.60 88.07 $\pm$ 2.43 91.22 $\pm$ 2.21 86.84 $\pm$ 1.91
VSSNet [35] 66.79 $\pm$ 0.62 60.87 $\pm$ 3.14 79.35 $\pm$ 0.48 67.91 $\pm$ 2.29 87.39 $\pm$ 1.22 65.90 $\pm$ 3.81 73.84 $\pm$ 1.17 73.14 $\pm$ 0.77 93.07 $\pm$ 0.58 74.95 $\pm$ 1.90
DSANet 78.09 $\pm$ 0.20 88.22 $\pm$ 1.52 87.32 $\pm$ 0.13 92.26 $\pm$ 1.56 86.40 $\pm$ 0.39 91.37 $\pm$ 1.57 88.59 $\pm$ 0.53 94.01 $\pm$ 1.40 92.98 $\pm$ 0.18 94.98 $\pm$ 1.39
  • ’*’ represents that the model input is a sequence and its MinIP.
TABLE IV: Quantitative comparisons with DSA segmentation and general segmentation methods on DSCA for the entire vessel. The results are average±$\pm$±standard deviation obtained from five-fold cross-validation. All metrics are based on percentages (%).
@@
Sequence + MinIP nnUNet* [22] 88.25 $\pm$ 0.15 82.34 $\pm$ 0.17 90.07 $\pm$ 0.01 88.65 $\pm$ 0.32 91.72 $\pm$ 0.31 94.12 $\pm$ 0.15
UNet* [8] 78.56 $\pm$ 1.13 71.70 $\pm$ 1.24 83.17 $\pm$ 0.86 88.38 $\pm$ 1.73 79.76 $\pm$ 2.26 93.58 $\pm$ 0.79
MISSFormer* [43] 80.95 $\pm$ 1.11 73.29 $\pm$ 1.34 84.34 $\pm$ 0.90 90.66 $\pm$ 0.64 79.54 $\pm$ 1.82 94.71 $\pm$ 0.27
H2Former* [44] 81.78 $\pm$ 0.39 74.05 $\pm$ 0.43 84.84 $\pm$ 0.30 87.89 $\pm$ 0.20 82.76 $\pm$ 0.43 93.46 $\pm$ 0.10
DeepMedic [30] 79.53 $\pm$ 0.89 73.51 $\pm$ 0.88 84.31 $\pm$ 0.64 85.20 $\pm$ 3.13 84.21 $\pm$ 3.43 92.20 $\pm$ 1.44
DSNet [34] 87.41 $\pm$ 0.10 81.65 $\pm$ 0.11 89.65 $\pm$ 0.07 88.24 $\pm$ 0.09 91.34 $\pm$ 0.19 93.91 $\pm$ 0.04
CAVE [45] 83.07 $\pm$ 1.88 75.89 $\pm$ 2.07 85.98 $\pm$ 1.44 87.05 $\pm$ 3.61 85.97 $\pm$ 1.64 93.15 $\pm$ 1.75
VSSNet [35] 82.72 $\pm$ 0.89 75.38 $\pm$ 0.90 85.67 $\pm$ 0.58 91.65 $\pm$ 0.72 81.37 $\pm$ 1.48 95.26 $\pm$ 0.31
@@
  • ’*’ represents that the model input is a sequence and its MinIP.
Figure 5: Segmentation results on the DSCA dataset. Green is MAT, and red is BV. Enlarged viewing for better clarity.
IV-C Cerebrovascular Segmentation Results
We extensively assessed our proposed method on the DSCA dataset, comparing it with state-of-the-art segmentation approaches, which include UNet [8], CENet [42], CSNet [16], AttentionUNet [17], SwinUNet [20], MISSFormer [43], H2Former [44], TransUNet [19], nnUNet [22], ERNet [31]. Additionally, we also reproduced [28] and [7] for comparison, as their codes are not publicly available. The comparative results of segmentation are shown in Table III.
We can observe that DSANet achieves the best performance across all metrics for segmenting both MAT and BV, which indicates the superiority of our method. In MAT segmentation, DSANet holds a significant advantage over other methods in terms of JAC, Dice, SEN, and AUC, surpassing H2Former by 4.32%, 3.92%, 4.50%, and 3.39%, respectively. A statistical significance of (p$<$0.05) was obtained using a paired t-test between the two methods. In BV segmentation, the pixels of BV constitute a relatively small proportion of the total image pixels, which explains why our proposed method is not significantly ahead of the comparative method. Nevertheless, DSANet surpasses nnUNet by 1.33% in JAC and 1.11% in SEN with statistical significance (p$<$0.05). Similarly, it also outperforms nnUNet by 0.87% in Dice and 0.57% in AUC, highlighting its superior performance with statistical significance (p$<$0.05).
For a more intuitive comparison of CA segmentation performance across different methods, we also provide a visual comparison in Fig. 5. As we can see the shape and background of the CA are very intricate, with low contrast in small vessels. In this challenging context, our DSANet effectively segments the CA, particularly ensuring small vessels with more robust connectivity, as indicated in the detailed orange boxes in ICA, ECA, and VA. In addition, skulls with deep contrast are mistakenly extracted as vessels in other methods, while our method successfully suppresses such false positives, as shown by the white arrows in ECA and VA in Fig. 5. In ICA, CENet and TransUNet perform well in segmenting vessels within the left orange box, while their performance diminishes in the right orange box. In contrast, DSANet achieves superior segmentation performance in both boxes. Furthermore, other methods erroneously segment the MAT in ECA (depicted in green), while DSANet effectively addresses this issue.
We also compare the performance of the entire CA without distinguishing BV and MAT to demonstrate the superiority of DSANet further. As indicated in Table IV, DSANet holds a slight advantage over other methods in JAC, Dice, SEN, and AUC, surpassing the second-place method by 1.06%, 0.64%, 0.85%, and 0.43%, respectively, with statistical significance (p$<$0.05). In PRE, AttentionUNet outperforms our method, but it falls behind our method in SEN. This indicates that AttentionUNet may miss some positive samples, reducing sensitivity, whereas our method achieves a better balance between SEN and PRE. To achieve precise segmentation, it is essential to focus on the intricate details of small vessel structures, which represent a minority of the total image pixels. Thus, clDice [41] is employed to assess vessel connectivity. DSANet significantly outperforms other methods, surpassing nnUNet by 1.04% in clDice with statistical significance (p$<$0.001).
Furthermore, for a fairer comparison, we have added experimental comparisons with model inputs as a sequence and its corresponding MinIP image in Table III and IV. These comparisons include nnUNet [22], UNet [8], MISSFormer [43], H2Former [44], DeepMedic [30], VSSNet [35], CAVE [45] and our previous work, DSNet [34]. Interestingly, for models like UNet, MISSFormer, and H2Former, their performance on metrics such as DICE was actually worse when using sequence input compared to just using the MinIP images. On the other hand, from the SEN metric, all of these methods saw a significant improvement, indicating that the introduction of the sequence increased the recall rate of the network, though this may also result in a decrease in precision. This could be because sequences contain more redundant information and background noise, making it crucial to handle these effectively. nnUNet can handle this issue effectively with its excellent training capabilities. nnUNet is a general model that outperforms some models specifically designed for DSA images, demonstrating its superior performance and generalization capabilities. However, as shown in Table IV, VSSNet surpasses nnUNet in SEN and AUC, despite being behind in other metrics. Additionally, compared to the two methods specifically designed for training with DSA sequences, CAVE and VSSNet, our DSANet significantly outperforms them in JAC, Dice, and PRE as shown in Table III and IV. Furthermore, it is noteworthy that our method performs slightly lower than VSSNet in terms of SEN and AUC metrics as shown in Table IV, which indicates that VSSNet tends to prioritize recall while neglecting precision. This result suggests that VSSNet is more inclined to capture true positives, while it may also yield a higher rate of false positives. Our method used nnUNet as the baseline and introduced TF and STF components, which can effectively learn spatio-temporal information, thereby achieving better performance.
TABLE V: Ablation study of these components of the proposed DSANet. The results are average±$\pm$±standard deviation obtained from the five-fold cross-validation experiment. All metrics are expressed as percentages (%).